Sliding on the Ice In the winter sport of curling, players give a 20 kg stone a push across a sheet of ice. The Slone moves approximately 40 m before coming to rest. The final position of the stone, in principle, onlyndepends on the initial speed at which it is launched and the force of friction between the ice and the stone, but team members can use brooms to sweep the ice in front of the stone to adjust its speed and trajectory a bit; they must do this without touching the stone. Judicious sweeping can lengthen the travel of the stone by 3 m. 80. A curler pushes a stone to a speed of 3.0 m/s over a time of 2.0 s. Ignoring the force of friction, how much force must the curler apply to the stone to bring it op to speed? A. 3.0 N B. 15 N C. 30 N <---CORRECT ANSWER D. 150 N 8 1. The sweepers in a curling competition adjust the trajectory of the slope by A. Decreasing the coefficient of friction between the stone and the ice.<---CORRECT ANSWER B. Increasing the coefficient of friction between the stone and the ice. C. Changing friction from kinetic to static. D. Changing friction from static to kinetic. 82. Suppose the stone is launched with a speed of 3 m/s and travel s 40 m before coming to rest. What is the approximate magnitude of the friction force on the stone? A. 0 N B. 2 N C. 20 N D. 200 N 83. Suppose the stone's mass is increased to 40 kg, but it is launched at the same 3 m/s. Which one of the following is true? A. The stone would now travel a longer distance before coming to rest. B. The stone would now travel a shorter distance before coming to rest. C. The coefficient of friction would now be greater. D. The force of friction would now be greater. ============================================================================================== I got the first two correct and the last two incorrect and I need some help. Please show calculations! Explanation would be nice as well. THANKS IN ADVANCE! Will rate!
Answer #1
D. 150 N

A. Decreasing the coefficient of friction between the stone and the ice.

B. 2 N
C. The coefficient of friction would now be greater
Answer #2
To find force, you need to know acceleration. To find that, you can use the final velocity equation:

Vfinal = Vinitial + a*t
2.9 = 0 + 2.1a
a = 1.381

F = ma = 1.381*20 = 27.62 N

Average power is the value of the total work done divided by the time it took to do the work.

Work = Force*Distance, so we need distance. We can use the displacement equation.

x = Vinitial*t + (1/2) at^2
x = 0 + (1/2) (1.381)(2.1)^2
x = 3.045 m

Work = F*d = 27.62*3.045 = 84.1 J

Average Power = W / t = 84.1 / 2.1 = 40 watts